OpenStudy (anonymous):

an isosceles triangular prism has a height of 9 in a base length of 7.2 and a base height of 6 in and base legs of 7 in what does it look like?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

thank you :D

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

and whats the big B ?

OpenStudy (mathstudent55):

Where's big B?

OpenStudy (anonymous):

the Big b is the area of the base \[B =bh\]

OpenStudy (mathstudent55):

The area of the base is A = (1/2)bh, where b = base of base triangle called here base length) and h = height of the base.

OpenStudy (anonymous):

im confused (/.\) homework blows

OpenStudy (anonymous):

how do i find the lateral surface area or this figure?

OpenStudy (mathstudent55):

The lateral surface is the total area of the three vertical sides. They are all rectangles. The front one has measures 7.2 in. and 9 in. The two side ones have the same measures of 7 in. by 9 in. You need to add the area of three rectangles, but two of them are the same. Lateral Area = (7.2 in. x 9 in.) + (7 in. x 9 in.) + (7 in. x 9 in.) Lateral Area = (64.8 + 63 + 63) in.^2 = 190.8 in.^2

OpenStudy (anonymous):

thank you alottt :D

OpenStudy (mathstudent55):

wlcm